OpenStudy (anonymous):

How do you figure out what the simplified form of things are? An example is sqrt 98n^19....

OpenStudy (dumbcow):

\[\sqrt{98n^{19}} = \sqrt{2*7^{2}*n^{18}n} = 7n^{9}\sqrt{2n}\]

OpenStudy (dumbcow):

one trick for variables inside square root, just divide the power by 2 only works if power is even, thats why i wrote n^19 as n^18*n

OpenStudy (anonymous):

Thanks! :)